Let S = {α, β, γ}, T = {α, α + β, α + β + γ}, W = {α + β, β + γ, α + γ} be subsets in (4)

a vector space V. Prove that L(S ) = L(T) = L(W).


1
Expert's answer
2021-07-14T14:52:31-0400

"s=(\alpha,\beta,\gamma)"

let "t \in L(T)" where t is a linear combination of element of T.

i.e. t="a(\alpha)+b(\alpha+\beta)+c(\alpha+\beta+\gamma)"

so "t\in" L(S)

L(T)"\subset" L(S).................................(1)

Now, w"\in" L(W) such that "w=a_1(\alpha+\beta)+b_1(\beta+\gamma)+c_1(\alpha+\gamma)"

w=(a1+c1)"\alpha"+

(a1+b1) "\beta" +(b1+c1)"\gamma"

w is a linear combination of "\alpha,\beta,\gamma"

so w"\in" L(S)................................(2)

now let s"\in" L(S), where "s=a_2+b_2\beta +c_3\gamma"

"s=a_1\alpha+b_2\beta+c_3\gamma"

we can write s in form of

"s=a_3\alpha +a_4(\alpha+\beta)+a_5(\gamma+\alpha)"

and

"s=a_6
(\alpha+\beta)+a_7(\beta+\gamma)+a_8(\gamma+\alpha)"

so, s"\in" L(T); s"\in" L(W)

"\therefore L(S)\subset L(T);
L(S)\subset L(W)"

"\therefore L(S)=L(T)=L(W)"

hence proved...